a balloon is floating around outside your window the temperature outside is 39 degree and the air pressure is 0.700 atm.your neighbor ,who released the balloon tell you that he filled it with it with 4.90 moles of gas what is the volume of gas inside this balloon
Answer #1

Given:

P = 0.7 atm

n = 4.90 mol

T = 39.0 oC

= (39.0+273) K

= 312 K

use:

P * V = n*R*T

0.7 atm * V = 4.9 mol* 0.08206 atm.L/mol.K * 312 K

V = 179.219 L

Answer: 179 L